# Range Membership Inference Attacks

### v2

@@ 4 Range membership inference attacks
4.2 Composite hypothesis testing
Similar to likelihood ratio tests for membership inference games 4.1, we can also construct two hypotheses for range membership inference game (Def 2):
@@ 4.2 Composite hypothesis testing
Since it can be intractable to iterate over all points in a given range $\mathcal{R}_{b}$, we use a set of sampled points $S$ in the given range as its proxy. In this way, the two hypotheses are:
[…]
The likelihood ratio in this case is still $\frac{\text{I\kern-1.04996ptP}{\left(\theta|H_{1}\right)}}{\text{I\kern-1.04996ptP}{\left(\theta|H_{0}\right)}}$. However, the alternative hypothesis $H_{1}$ is composite because it is a union of multiple hypotheses $\bigcup_{z_{i}\leftarrow S}(z_{i}\in D)$. Therefore, we need to use statistical methods tailored for composite hypothesis testing. There are two commonly used methods for it: Bayes Factor [22] and Generalized Likelihood Ratio Tests (GLRTs) [43].
@@
GLRT assumes that the true hypothesis $h^{*}$ is explicitly listed in the composite hypothesis, which, in our case, means the training point is in the set of samples $\{z_{i}\}$. In this case, the composite hypothesis can be reduced to a single hypothesis by taking the maximum:
@@
Bayes Factor assumes the hypothesis is a random variable, each with a likelihood value according to a prior distribution. In our case, it means each point $z$ in the range can be sampled with probability determined by the prior data distribution $\text{I\kern-1.49994ptP}{(z)}$. To reduce the composite hypothesis to a simple one, Bayes Factor computes the weighted average:
@@
This can be thought of as taking the expectation of the composite hypothesis or using the most representative hypothesis instead of the maximal (as in GLRT) in the likelihood ratio test.
@@
At first look, GLRT might be an intuitive choice, because taking the maximum leads to a two-step strategy: search and test. Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points, and their membership should be indicative of the ranges’ membership. However, this assumes the true training point is included in every sampled set created by the adversary, which is extremely unlikely. Secondly, this also assumes that we can reliably find the max values in a given range. Since most ranges are large data subspaces, it is very challenging to find the optimal points within the large space. Even if the search space can be navigated, any search algorithm is likely to return local maxima.
To make full use of the Bayes Factor, we need to know the prior distribution, which is also not possible.
A more important issue is that the power of existing statistical methods relies on the absolute correctness of the likelihood values: the larger the likelihood values, the more likely the point is a member. In other words, the largest likelihood values of non-members must be smaller than the smallest likelihood of members. However, membership inference attacks are known to be non-perfect; they are also known to be particularly unreliable on out-of-distribution (OOD) data [48] and can assign them high scores. When the sampling space contains only these data as opposed to real and in-distribution (ID) data, the maximum might not be anything close to true training data, increasing the FPR and lowering the AUC as a result.
4.3 Our approach: Trimmed averages
Given these limitations, in this paper, we propose a simple attack strategy based on Bayes Factor. We first assume every sampled point is of equal probability. Then, to address the unreliability of MIA scores, we use a trimming process based on the type of data in the sampling space. If the data are all naturally ID, we trim the less probable samples and take the average likelihood of the top samples to reduce the influence of non-members and the randomness in the sampling process. On the other hand, if the adversary can only synthesize data within the range instead of sampling from a pool of real data, the top samples are highly likely to be OOD data with high membership scores. Hence, in this case, we want to remove those points from the equation. We then trim the top-scoring samples and take the average of the rest. This gives us the following:
@@ 4.3 Our approach: Trimmed averages
where $S$ is the sampled set, $q_{s}$ and $q_{e}$ mark the start and the end of the quantiles that we want to remove to compute our robust statistics that are one-sided trimmed means. If the sampling space is filled with synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points. $q_{s}$ is a hyperparameter that decreases (trim more) as the quality of sampled points gets worse. On the other hand, if the sampling space consists of real points, we set $q_{s}=0$ to remove the smallest points in our aggregation. $q_{e}$ decreases (trim less) as the number of real samples decreases to offset the high variance due to limited samples available. If we do not trim anything, the formulation reduces to Bayes Factor which assigns equal probability to all points.
Note that the optimal hyperparameters may differ across different membership signals (e.g. loss values, LiRA scores), as they exploit different vulnerabilities and expose different training points. However, for fixed model architectures, range functions, data distributions, and sampling methods, these hyperparameters can be determined by reference models, similar to the offline version of RMIA [48]. Specifically, by randomly choosing a reference model as the temporary target model, we can run RaMIAs using all other reference models while sweeping these hyperparameters in a grid search manner. More details are in Appendix A.
4.4 Range membership inference attack as a framework
The range membership inference attack is a new inference attack framework, not a particular attack algorithm. There are two components in this framework: a sampler and a membership tester, both of which are necessary to compute the range membership score formulated in Eqn 4. The sampler $\text{Sampler}(\mathcal{R}):S\rightarrow X$ returns samples within the given range. The membership tester $\text{MIA}(x)$ is a (point-query) membership inference algorithm that outputs a membership score, which can be used to approximate $\text{I\kern-1.49994ptP}{(\theta|x)}$. Any existing MIA algorithm can be plugged in as RaMIA’s membership testing backbone. Similar to MIAs, the key to using RaMIA as a privacy auditing tool is to compute the range membership score. Our framework can adopt any existing membership scoring function $\text{MIA}(x)$ to compute $\text{RaMIA}(\mathcal{R})$. Below, we outline the attack with our attack strategy described above:
@@ 4.4 Range membership inference attack as a framework
5 Experiments
Since the purpose of this paper is to introduce a new concept and framework, the goal of the experiments section is to provide a proof-of-concept. We experiment on the commonly used Purchase-100 [40], CelebA [28], CIFAR-10 [25] and AG News [49] datasets. We will describe the details of how we split the dataset, train models, construct ranges, and obtain samples in Section 5.1. Since the range membership notion is new, we do not have a prior method to compare with. However, since the goal of introducing this new privacy notion is to enable better and more comprehensive privacy auditing, we compare RaMIA with the de-factor privacy auditing framework, MIA. We show that RaMIA performs better (Sec 5.5) in scenarios where MIA under-performs (Depicted in Figure 2). Both MIA and RaMIA in the experiments are built upon the state-of-the-art attack algorithm, robust membership inference attack (RMIA) [48], with a few reference models trained in the same way as [5] and [48]. The respective queries are outlined in Table 1, and the definitions of members under each attack framework are explained in Table 2. In both tables, $x$ represents original data in datasets, while $x^{\prime}$s are either data with missing values or modified data from $x$. The reason that we do not test our attacks by taking ranges centered at original data $x$ is that the chance of the attack data being exactly the same as the training data is extremely low without sufficient prior knowledge. It is more realistic that similar data are being queried.
Table 1: Range queries and point queries used in our experiments for RaMIA and MIA respectively.
@@ 5 Experiments
Table 2: Definitions of range and point members corresponding to the attack queries in Table 1.
@@
5.1 Setup details
As mentioned earlier, the range function must come from the semantics of privacy. Hence, in experiments, we use specific range functions with different data types. For tabular data, we consider missing columns, which is an extreme case of using Euclidean distance on missing columns as a range function. For human photo data, we use a semantic range function, which is based on the main person featured in the photo. For other image data, we use geometric transformations as range functions. For text data, we use (word level) Hamming distance, which is edit distance or Levenshtein distance that only considers word substitution. The reasons for choosing these range functions have been motivated and explained in earlier parts of this paper.
5.1.1 Tabular data: Purchase-100
  • Dataset Purchase-100 [40] is a tabular dataset derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data. This dataset was first curated by [40] such that there are 600 binary features, each representing if each person, represented by each row, has purchased the product. The data is then divided into 100 classes, and the task is to predict the category of the person given the purchase history.
  • Models We train five-layer multi-layer perceptron (MLP) models in PyTorch [34] on half of the entire dataset. The hidden layers are of sizes $[512,256,128,64]$, identical to the setup in [48]. All models achieve a test accuracy of $83\%$.
  • Construction of ranges We simulate the scenario where the attacker has incomplete data (data with missing values). For all training and test data records, we randomly mask k columns. Each row with masked columns is a range query that contains $2^{k}$ possible points as each feature is binary. We then check if each range constructed from test data points includes any training point. If so, they are re-labeled as ”IN-ranges”.
  • Sampling within ranges Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns. The parameter of the sampler is computed by taking the average value of each column. Because of the nature of this dataset, our sampled data can be regarded as in-distribution. We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
5.1.2 Image data I: CelebA
  • Dataset CelebA [28], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features. We construct the members set by only including photos of celebrities with identity numbers smaller than 5090. The rest are used to construct the non-members set. For each celebrity in the members set, half of the photos are put into the training set, while the other half goes into the holdout set.
@@ 5.1.2 Image data I: CelebA
  • Construction of ranges The range function here is a semantic one that is based on the identity of the face image. For example, a range query can be ”all Alice’s photos”. Since the identities in the training and non-members set are disjoint, it is easy to construct IN- and OUT-ranges based on the distribution of identities in the two sets.
  • Sampling within ranges For each range query, we curate all images in the holdout set that share the same identity as the range center to construct our sample set.
5.1.3 Image data II: CIFAR-10
  • Dataset CIFAR-10 [25] is a popular image classification dataset. There are 50,000 training images, each of size $(32,32,3)$.
  • Models We train WideResNets-28-2 [47] with JAX [2] on half of the training set of CIFAR-10 using the code from [5], with and without image augmentations. Our target model trained without augmentation achieves a test accuracy of $83\%$ on the CIFAR-10 test set, and the target model trained with augmentation achieves a test accuracy of $92\%$. The train time augmentation is the composition of random flipping, cropping and random hue.
  • Construction of ranges The range function here is different types of image augmentations, which are geometric transformations. An example of a range query is ”all transformed version of image $X$”. For each training and test image, a range is constructed by applying different transformations.
  • Sampling within ranges For each range query, we independently apply 15 image augmentations on the query center. The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
@@ 5.1.3 Image data II: CIFAR-10
Figure 3: RaMIAs perform better than MIAs on points that are close to original points but not exactly the same.
5.1.4 Textual data: AG News
@@
  • Dataset We use the AG News dataset [49], which is a news collection with four categories of news. Although it was introduced as a text classification dataset, we disregard the labels and treat it as a text generation dataset. There are 120,000 sentences in its training set. It is a popular text dataset in LLM MIA literature [33].
@@
  • Models We took pretrained GPT-2 [35] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [19] implemented in Hugging Face’s PEFT [32] library. The finetuning is done for 4 epochs. Our target model achieves a perplexity of 1.39 on the test set of AG News.
@@
  • Construction of ranges The range function here is word-level Hamming distance, which can be thought of as the edit distance measured on word level that only allows word substitution. An example of a range query is ”all sentences within Hamming distance $d$ to sentence $x$”. To construct IN- and OUT-ranges, we just need to specify the max Hamming distance and the starting sentence. We constructed the starting sentences by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pretrained BERT [12] model, so they have a distance of $\alpha$ to the original training/test sentences. A Hamming distance is then specified with each starting sentence to form a range.
@@
  • Sampling within ranges We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [12] to replace the mask with one of the top choices.
5.2 Hyperparameters
@@
Overall, on Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$. On CIFAR-10, we apply up to 15 distinct transforms, and set $q_{e}=100,q_{s}=40$. On AG News, we construct 50 sentences within each range, and set $q_{e}=100,q_{s}=20$. On CelebA, each celebrity has a different number of images in the sampling space, ranging from 1 to 18. Since it is hard to standardize the sample size for all ranges, we take all of them. We then set $q_{s}=0$ and $q_{e}=25$, which means we are not trimming anything for ranges with very few samples available.
5.3 Implementation details

### v3

@@ IV Range membership inference attacks
IV-B Composite hypothesis testing
When extending to range membership inference, the adversary is presented with a set of points rather than a single point. For a given range $\mathcal{R}_{b}$, we define the hypotheses as:
@@ IV-B Composite hypothesis testing
Since it can be intractable to iterate over all points in a given range $\mathcal{R}_{b}$, we use a set of sampled points $S$ in the given range as its proxy to replace $\mathcal{R}_{b}$ in the hypotheses above.
The likelihood ratio in this case is still $\frac{\text{I\kern-1.04996ptP}{\left(\theta|H_{1}\right)}}{\text{I\kern-1.04996ptP}{\left(\theta|H_{0}\right)}}$. However, the alternative hypothesis $H_{1}$ is composite, representing a union of multiple hypotheses $\bigcup_{z_{i}\leftarrow S}(z_{i}\in D)$. Therefore, we need to use statistical methods tailored for composite hypothesis testing. Two commonly used methods are Bayes Factor [31] and Generalized Likelihood Ratio Tests (GLRTs) [32].
@@
GLRT assumes that the true hypothesis $h^{*}$ is explicitly present in the composite hypothesis, which, in our case, means the training point is in $S$. This allows us to reduce the composite hypothesis by taking the maximum over $S$:
@@
Bayes Factor treats the hypothesis as a random variable with a prior distribution. Each point z in the range is sampled with probability determined by $\text{I\kern-1.49994ptP}{(z)}$, and the composite likelihood is approximated via the expectation:
@@
At first look, GLRT might be an intuitive choice, as it is equivalent to a two-step strategy: search and test. Searching for the points with the highest membership score is conceptually equivalent to identifying the points that are most likely to be training points, and their membership should be indicative of the ranges’ membership. However, this assumes the true training point is in every sampled set $S$, which is extremely unlikely. Secondly, this also assumes that we can reliably find the max values in a given range. Since most ranges are large data subspaces, it is very challenging to find the extreme points. Even if the search space can be navigated, search algorithms are likely to return local maxima.
Similarly, the Bayes Factor approach requires knowledge of the true prior $\text{I\kern-1.49994ptP}{(x)}$, which is generally unknown. Both methods also hinge on the absolute correctness of the likelihood values: ideally, the highest likelihood among non-members should be lower than the lowest likelihood among members. However, membership inference attacks are known to be imperfect and particularly unreliable on OOD data [18], which may assign high scores erroneously to non-members. This leads to increased Range FPR and reduced AUC, especially when working with sampled data.
IV-C Our approach: Trimmed averages
To overcome these limitations, we propose a robust attack strategy based on a modified Bayes Factor approach that employs trimmed averages. Our method begins by assuming every sampled point in $S$ is equally probable. To mitigate the unreliability of MIA scores, we introduce a trimming process that adapts to the nature of the data in the sampling space:
  • In-Distribution (ID) Data: When the sampled points are naturally in-distribution, we trim the lower quantiles (i.e., those with the smallest likelihood values) and average the top samples. This reduces the influence of non-members and random noise.
  • Synthetic Data: When only synthetic data are available, the highest scoring samples are often OOD and prone to false positives. In this case, we trim the top quantiles and average the remaining samples.
This strategy is formalized as:
@@ IV-C Our approach: Trimmed averages
where $S$ is the sampled set, $q_{s}$ and $q_{e}$ denote the quantile thresholds between which to be trimmed. For synthetic data, the chance of the top samples being false positives is high, so we set $q_{e}=100$ to remove the largest points. $q_{s}$ is a hyperparameter that decreases (trim more) as the quality of sampled points gets worse. For real data, we set $q_{s}=0$ to remove the smallest points in our aggregation. $q_{e}$ decreases (trim less) as the number of real samples decreases to offset the high variance due to limited samples available. If we do not trim anything, the formulation reduces to Bayes Factor which assigns equal probability to all points. We provide an empirical comparison in Fig 11 to show that the trimmed means are better suited for RaMIA.
Note that the optimal hyperparameters may vary with different membership signals (e.g., loss values, LiRA scores) since these signals capture different vulnerabilities. For fixed model architectures, range functions, data distributions, and sampling methods, these hyperparameters can be determined using reference models, similar to the offline version of RMIA [18]. By randomly choosing a reference model as a temporary target model and sweeping through hyperparameters via grid search, one can identify the best values (details in Appendix -A).
IV-D Range membership inference attack as a framework
Our proposed RaMIA is not a single attack algorithm but a new inference attack framework. It comprises two key components:
  • Sampler: $\text{Sampler}(\mathcal{R}):\mathcal{R}\rightarrow S$ returns samples within the given range
  • Membership Tester: MIA(x) is any point-query membership inference algorithm that outputs a membership score, approximating $\text{I\kern-1.49994ptP}{(\theta\mid x)}$. This function can be replaced by any existing MIA algorithm.
The core idea is to compute a range membership score $\text{RaMIA}(\mathcal{R})$ by robustly aggregating the membership scores of the sampled points $S$ using our trimmed average approach (Eqn. 4). This framework leverages existing membership scoring functions while addressing the new challenge of capturing privacy leakage in a broader neighborhood.
@@ IV-D Range membership inference attack as a framework
Importantly, our formulation of RaMIA as a composite hypothesis test—and the subsequent development of the trimmed averages approach—represents a first, yet promising, attempt to tackle this new technical challenge. While there is room for further exploration and refinement, our approach demonstrates that extending the membership inference framework to encompass ranges can capture significant privacy leakage in the vicinity of training data. This represents a core contribution of our work and lays the groundwork for future research in comprehensive privacy auditing.
V Experiments
Since the purpose of this paper is to introduce a new concept and framework, the goal of the experiments section is to provide a proof-of-concept.
We experiment on the commonly used Purchase-100 [13], CelebA [33], CIFAR-10 [34] and AG News [35] datasets. Details on dataset splits, model training, range construction, and sample acquisition are described in Section V-A. Since the range membership notion is new, we do not have a prior method to compare with. However, as our aim in introducing this new privacy notion is to enable better and more comprehensive privacy auditing, we compare our RaMIA framework with the de facto privacy auditing framework, MIA. Note that It is crucial to emphasize that our experimental comparison between RaMIA and MIA is not based on the standard MIA notion of membership. In standard MIAs, only the exact training data are considered members. In our experiments, the queries for MIA are the range centers—which are not training data by definition. Consequently, if one were to evaluate MIA using the correct MIA privacy notion, the AUC would be close to random guessing (i.e., 0.5). Instead, we use the MIAScore of the range center to solve the range membership inference game, even though this deviates from the standard definition. This allows us to compare the power of both frameworks in identifying queries that leak privacy.
Table I outlines the range queries and point queries used for RaMIA and MIA respectively, while Table II defines the notion of members under each attack setting. In both tables, $x$ represents the original data in the datasets, while $x^{\prime}$ represents either data with missing values or modified data derived from $x$. The reason we do not center ranges at the original data $x$ is that, without sufficient prior knowledge, the probability of the attack data exactly matching a training point is extremely low. In practice, similar but not identical data are more likely to be queried. It also acts as a hard case for RaMIA.
TABLE I: Range queries and point queries used in our experiments for RaMIA and MIA respectively.
@@ V Experiments
TABLE II: Definitions of range and point members corresponding to the attack queries in Table I.
@@
V-A Setup details
As mentioned earlier, the range function must derive from the semantics of privacy. Hence, in experiments, we use specific range functions tailored to different data types. For tabular data, we consider missing columns, which is an extreme case of using Euclidean distance on missing columns as a range function ($R=\{x^{\prime}:d(x^{\prime}_{i},x_{i})=0\land d(x^{\prime}_{j},x_{j})<C_{j}\}$, where $j$’s are masked columns and $i$’s are observable columns, $d$ is Euclidean distance, $C$ is an upper bound for each missing column to make sure $x^{\prime}_{j}$ is bounded by the infinity or extreme values (e.g., Age.). For human photos, we use a semantic range function based on the main person featured. For other image data, we use geometric transformations as range functions. For text data, we use (word level) Hamming distance, which is edit or Levenshtein distance that only considers word substitution. The reasons for choosing these range functions have been motivated and explained in earlier parts of this paper.
V-A1 Tabular data: Purchase-100
  • Dataset Purchase-100 [13] is derived from Kaggle’s Acquire Valued Shoppers Challenge  [^1] https://www.kaggle.com/c/acquire-valued-shoppers-challenge/data. It contains 600 binary features, representing the purchase history of each person. The data is divided into 100 classes. The task is to predict a person’s category given the purchase history.
  • Models We train a four-layer multi-layer perceptron (MLP) in PyTorch [36] on half of the entire dataset. The hidden layers are of sizes $[1024,512,256]$. All models achieve a test accuracy of $86\%$.
  • Construction of ranges We simulate the scenario where the attacker has incomplete data (data with missing values). For all training and test data records, we randomly mask k columns. Each row with masked columns is a range query that contains $2^{k}$ possible points. We re-labeled any range that contains at least one training point as ”IN”.
  • Sampling within ranges Since this dataset contains 600 independent binary features, we do Bernoulli sampling independently for all missing columns. The parameter of the sampler is computed by taking the average value of each column. Due to nature of this dataset, our sampled data can be regarded as ID. We take 19 samples for each range, together with the data obtained by doing mode imputation (fill in the missing values with the modes).
V-A2 Image data I: CelebA
  • Dataset CelebA [33], also known as the CelebFaces Attributes dataset, contains 202,599 face images from 10,177 celebrities, each annotated with 40 binary facial features. We construct the members set by only including photos of the first 5090 celebrities. The rest are used to construct the non-members set. For each celebrity in the members set, half of the photos are put into the training set, while the other half goes into the holdout set.
@@ V-A2 Image data I: CelebA
  • Construction of ranges The range function here is semantic, defined by identity. For example, a range query can be ”all photos of Alice”. Since the identities in the training and non-members set are disjoint, it is easy to construct IN and OUT ranges.
  • Sampling within ranges For each range query, we collect all holdout images sharing the same identity as the range center.
V-A3 Image data II: CIFAR-10
  • Dataset CIFAR-10 [34] is a popular image classification dataset, containing 50,000 training images of size $(32,32,3)$.
  • Models We train WideResNets-28-2 [37] with JAX [38] on half of the training set of CIFAR-10 using the code from [17], with and without image augmentations. Our target model achieves a test accuracy of $83\%$ when trained without augmentation, and $92\%$ with augmentation. The train-time augmentation is the composition of random flipping, cropping and random hue.
  • Construction of ranges The range function is defined by a set of geometric transformations (e.g., flipping, rotation, cropping). A range query consists of various transformed versions of image $x^{\prime}$.
  • Sampling within ranges For each range query, we independently apply 10 image augmentations to the range center. The augmentations include flipping, random rotation, random resizing and cropping, random contrast, brightness, hue, and the composition of them.
@@ V-A3 Image data II: CIFAR-10
Figure 4: RaMIAs perform better than MIAs on similar but not identical data points to training points that carry overlapping private information. This is the hardest setting for RaMIA where there is only one training point in each IN range, and they are not situtated at the range centers. We will show how RaMIAs will perform better in less restricted scenarios in a later figure. Note that the TPRs/FPRs are the range versions, and the MIA is the approach that uses the range center’s MIA score as the range’s membership signal, both introduced in Sec V-D.
V-A4 Textual data: AG News
@@
  • Dataset We use the popular AG News dataset [35], which is a news collection with four categories of news. We treat it as a text generation dataset, disregarding their labels. It contains 120,000 sentences in the training set.
@@
  • Models We took pretrained GPT-2 [39] models from Hugging Face’s transformers library, and finetuned them on half of AG News’ training set with LoRA [40] (implemented in Hugging Face’s PEFT [41] library). The finetuning is done for 4 epochs. Our target model achieves a test perplexity of 1.39.
@@
  • Construction of ranges The range function here is word-level Hamming distance, which can be thought of as the edit distance measured on word level that only allows word substitution. An example of a range query is ”all sentences within Hamming distance $d$ to sentence $x$”. To construct IN and OUT ranges, we construct range centers by randomly masking $\alpha$ words from the training and test sentences, before filling in the mask with a pretrained BERT [42] model, so they have a distance of $\alpha$ to the original training/test sentences.
@@
  • Sampling within ranges We mask the range center by $k$ words where $k$ is the Hamming distance specified by the range. Then we use BERT [42] to complete the masks.
V-B Hyperparameters
@@
Overall, on Purchase-100, we take 20 samples in every range, and set $q_{e}=100,q_{s}=45$. On CIFAR-10, we apply up to 10 distinct transforms, and set $q_{e}=100,q_{s}=40$. On AG News, we construct 50 sentences within each range, and set $q_{e}=100,q_{s}=20$. On CelebA, each celebrity has a different number of images in the sampling space, ranging from 1 to 18. Since it is hard to standardize the sample size for all ranges, we take all of them. We then set $q_{s}=0$ and $q_{e}=25$, which means we are not trimming anything for ranges with very few samples available.
V-C Implementation details